2. Your phone slips from your pocket as you’re going up a ladder of a high building. Your phone falls for 2.7 seconds before you heart it hit the cement below.






a. At what height did the phone fall from? (Hint: Your phone’s initial velocity is zero.)






b. What was the velocity of your phone when it hit the cement?


1
Expert's answer
2021-10-13T15:41:05-0400

a)


"h=0.5(9.8)(2.7)^2=35.7\ m"

b)


"v=(9.8)(2.7)=26.5\frac{m}{s}"